OpenStudy (anonymous):

Leslie wrote this equation for the area of a parallelogram. 17.8h = 384.48. What is the height (h) of this parallelogram?

Mehek (mehek14):

divide \(\dfrac{384.48}{17.8}\)

OpenStudy (anonymous):

thx

Mehek (mehek14):

yw :) and Welcome to OS

OpenStudy (anonymous):

can you help with a few more? @Mehek14

Mehek (mehek14):

sure

OpenStudy (anonymous):

k hang on

OpenStudy (anonymous):

The length of one side of a square is 58 inches. What is the area of the square? A. 29 in2 B. 3364 in2 C. 58 in2 D. 7.6 in2

Mehek (mehek14):

58 * 58 = ?

OpenStudy (anonymous):

3364

Mehek (mehek14):

ye correct

OpenStudy (anonymous):

thx

Mehek (mehek14):

yw

OpenStudy (anonymous):

1 sec i have more

OpenStudy (anonymous):

solve for n (10 5/8 + 3 5/8) 2 3/8 = n + 12

Mehek (mehek14):

the equation is not so clear :/

OpenStudy (anonymous):

i had to type it, eveerything is an image. 10 5/8 is a fraction. as is the other numbers like that

Mehek (mehek14):

oke i see the stuff in the parenthesis wut about outside? is it multiplication?

OpenStudy (anonymous):

+

Mehek (mehek14):

\((10\dfrac{5}{8}+3\dfrac{5}{8})+2\dfrac{3}{8}=n+12\) ?

OpenStudy (anonymous):

yep

Mehek (mehek14):

do you know how to add fractions?

OpenStudy (anonymous):

kinda... i suck balls at math

Mehek (mehek14):

so add everything on the left side

OpenStudy (anonymous):

?

Mehek (mehek14):

add the fractions \(10\dfrac{5}{8}+3\dfrac{5}{8}\)

OpenStudy (anonymous):

ok

OpenStudy (mathmale):

In this case we can simply throw out the parentheses on the left side. Add any two of the mixed numbers together, and then add to this sum the third number.

OpenStudy (anonymous):

14 1/4

Mehek (mehek14):

corect now add the other fraction

OpenStudy (anonymous):

4 5/8? @Mehek14

Mehek (mehek14):

no

OpenStudy (anonymous):

:(

Mehek (mehek14):

\(14\dfrac{1}{4}+2\dfrac{3}{8}\)

OpenStudy (anonymous):

i mean is the answer 4 5/8

OpenStudy (anonymous):

to the whole question

Mehek (mehek14):

ohhh oke hold on

OpenStudy (kesumonu):

37/8

OpenStudy (anonymous):

thats not a choice even

Mehek (mehek14):

ye 4 5/8

OpenStudy (anonymous):

k thx

Mehek (mehek14):

yw :)

OpenStudy (anonymous):

What is the reciprocal of 2 5/7 ?

OpenStudy (kesumonu):

either u r joking or a joke sorry to offend

OpenStudy (anonymous):

what do you mean?

OpenStudy (kesumonu):

i mean 7/25, u r insulting maths

OpenStudy (anonymous):

i meant 2 and 5 over 7, like a fraction, and "u r insulting maths" id horrible grammar. you are a joke.

OpenStudy (kesumonu):

btw , this is maths , i dont know grammer,

OpenStudy (kesumonu):

may be

OpenStudy (kesumonu):

adieu

OpenStudy (anonymous):

its math not maths